OpenStudy (anonymous):

The shadow of a tree on the ground is 17 feet. If the shadow of a 5 1/2 feet tall person standing up straight is 11 feet, how tall is the tree? I know the answer is 8 1/2 ft. but how do you solve it?

OpenStudy (zpupster):

it is a ratio 11 is to 5 1/2 as 17 is to ???

OpenStudy (anonymous):

8 1/2

OpenStudy (zpupster):

yes that is right, but how did you solve it

OpenStudy (anonymous):

that's what I don't know how to do, can you help me please

OpenStudy (zpupster):

sure

OpenStudy (anonymous):

thanks

OpenStudy (zpupster):

I set them like this: 11/( 1/2) = 17/x and then solve for x like so 11 x = 17(5 1/2) x = 17(5 1/2)/11 do the math and you should come out with your answer I cross mult and then divided to isolate x

OpenStudy (anonymous):

oh that makes sense thank you! :)

OpenStudy (zpupster):

good job!!

OpenStudy (anonymous):

thanks you now I get it :D